# RS-MoE: Mixture of Experts for Remote Sensing Image Captioning and Visual Question Answering

### v1

@@ Abstract
Remote Sensing Image Captioning (RSIC) presents unique challenges and plays a critical role in applications such as environmental monitoring, urban planning, and disaster management. Traditional RSIC methods often struggle to produce rich and diverse descriptions. Recently, with significant advancements in Vision-Language Models (VLMs), efforts have emerged to integrate these models into the remote sensing domain and to introduce richly descriptive datasets specifically designed to enhance VLM training. However, most current RSIC models generally apply only fine-tuning to these datasets without developing models tailored to the unique characteristics of remote sensing imagery. This paper proposes RS-MoE, a first Mixture of Expert based VLM specifically customized for remote sensing domain. Unlike traditional MoE models, the core of RS-MoE is the MoE Block, which incorporates a novel Instruction Router and multiple lightweight Large Language Models (LLMs) as expert models. The Instruction Router is designed to generate specific prompts tailored for each corresponding LLM, guiding them to focus on distinct aspects of the RSIC task. This design not only allows each expert LLM to concentrate on a specific subset of the task, thereby enhancing the specificity and accuracy of the generated captions, but also improves the scalability of the model by facilitating parallel processing of sub-tasks. Additionally, we present a two-stage training strategy for tuning our RS-MoE model to prevent performance degradation due to sparsity. We fine-tuned our model on the RSICap dataset using our proposed training strategy. Experimental results on the RSICap dataset, along with evaluations on other traditional datasets where no additional fine-tuning was applied, demonstrate that our model achieves state-of-the-art performance in generating precise and contextually relevant captions. Notably, our RS-MoE-1B variant achieves performance comparable to 13B VLMs, demonstrating the efficiency of our model design. Moreover, our model demonstrates promising generalization capabilities by consistently achieving state-of-the-art performance on the Remote Sensing Visual Question Answering (RSVQA) task.
@@ I Introduction
Image captioning integrates computer vision and natural language processing to automatically generate descriptive text for images, effectively bridging the information gap between visual and linguistic domains. Unlike captioning of common natural images, remote sensing image captioning (RSIC) poses greater challenges due to the inherent characteristics of remote sensing images, which cover larger geographic areas and more diverse geographical objects. Consequently, RSIC tasks demand not only the description of geographic object information but also a comprehensive understanding of the relationships and interactions among various geographic objects.
Early studies of RSIC focused on traditional methodologies, including template-based  [1] and retrieval-based methods  [2, 3]. However, these methods cannot generate rich and varied descriptive sentences. Recent studies [4, 5, 6, 7] commonly adopt the encoder-decoder architecture, which divides the RSIC task into an image encoding phase that extracts semantic features from the input image, and a sequence modeling phase that uses the extracted features to generate text and sentences. Depending on the specific models employed, encoder-decoder approaches can be further categorized into CNN-based encoders with RNN/LSTM decoders, and CNN-based encoders with Transformer decoders. Although these methods have achieved satisfying performance on the RSIC task, they typically generate only one or two simple sentences for captioning, limiting their practical application. This limitation typically arises due to two principal reasons: the simplicity and repetitiveness of the sentences in the datasets used for training models; and the relatively limited ability of these models to extract semantic features and generate complex descriptions.
The Large Language Models (LLMs) and Vision Language Models (VLMs) have recently achieved significant success across multiple fields, including computer vision [8], natural language processing [9], and robotics [10]. Particularly, VLMs [11, 12, 13] have effectively narrowed the gap between visual images and textual natural language by advancing the understanding of intermodal relationships, reaching a level of visual comprehension comparable to human capabilities. Therefore, some researchers [14, 15, 16, 17, 18, 19, 20] have recently shifted their focus to applying VLMs to remote sensing visual interpretation tasks. Specifically, for RSIC tasks, the RSGPT [15] introduced the human-annotated RSICap dataset, which provides high-quality and detailed descriptions for remote sensing images. The remote sensing captions in this dataset include summarized scene descriptions, object information (color, shape, and count), and object relationships (relative position). However, RSGPT merely fine-tunes an existing VLM model on this dataset without proposing a novel model to explore and utilize the provided detailed descriptions deeply.
@@
  • We are the first to introduce the MoE framework within the domain of remote sensing. By drawing on the idea behind MoE for handling complex tasks through task decomposition, we break down the intricate challenges of remote sensing image captioning into distinct subtasks, enabling specialized expert models to more effectively address the diversity and complexity inherent in remote sensing data.
  • We propose a novel Vision-Language Model, RS-MoE, specifically designed for remote sensing image captioning. Unlike traditional MoE models, our model introduces two key innovations: a novel Instruction Router module that dynamically generates task-specific prompts and lightweight LLMs as expert models. These innovations optimize the model’s ability to handle the complexity and diversity of remote sensing images while maintaining computational efficiency.
  • We present a two-stage training strategy for the RS-MoE model tailored to the RSIC task, which incorporates a proper initialization process to prevent model degradation due to the sparsity that typically results from directly applying MoE. Additionally, we employ the LoRA technique to significantly reduce the number of trainable parameters, thereby enhancing the efficiency and manageability of the model during the training process.
  • Extensive experiments demonstrate that, even with fine-tuning on only one dataset containing around 3,000 images, our model achieves state-of-the-art performance across four RSIC datasets and exhibits strong generalization capabilities on an RSVQA dataset. Notably, our lightweight RS-MoE-1B model achieves comparable performance to larger 13B VLM methods for remote sensing image captioning, while offering significantly greater efficiency.
Figure 1: Overview of the proposed RS-MoE model, which consists of three key components: the Image Encoder, the VLM Encoder, and the MoE Block. The MoE Block comprises an Instruction Router that dynamically generates task-specific prompts and three lightweight LLMs, which focus on different aspects of the captioning task. In the generated captions, shown in the top right corner of the figure, distinct colors represent each aspect: orange for the overall theme, purple for specific objects, and green for relationships between objects. RS-MoE is trained using a novel two-stage training strategy specifically designed for remote sensing image captioning. In Stage I (a), the VLM Encoder and the LLM Block are fine-tuned to initialize model weights specifically designed for the RSIC task. In Stage II (b), the MoE Block is fine-tuned to produce more detailed captions for RSIC tasks.
@@ II-B VLM in Remote Sensing
Vision-Language Models (VLMs) [11, 12, 13] are cutting-edge frameworks designed to integrate visual and textual modalities, enabling models to effectively interpret and generate both visual content and corresponding linguistic descriptions. These models have demonstrated significant advancements in tasks involving natural images, especially in generating dialogues and interpreting visual data. By combining visual features with contextual cues, VLMs provide a more interactive and coherent understanding of visual inputs. For instance, the LLaVA model [21] exemplified both data efficiency and architectural simplicity by employing a fully connected vision-language projection layer, enabling multimodal understanding using only 600K image-text pairs. More recently, Qwen2-VL [22] introduced a dynamic resolution mechanism alongside Multimodal Rotary Position Embedding M-RoPE, which allows the model to process images of varying resolutions, substantially improving accuracy and flexibility in multimodal tasks.
@@
Inspired by the rapid advancements in computer vision, researchers have increasingly focused on integrating Vision-Language Models (VLMs) into remote sensing tasks. One of the early comprehensive reviews,  [23], highlighted key VLM applications within the remote sensing domain, identifying both the current challenges and future directions for research. Since then, several VLM-based models have emerged to address these challenges. For example,  [18] introduced EarthGPT, a model specifically designed for multi-sensor remote sensing image interpretation. This model integrates three core innovations: vision-enhanced perception, cross-modal understanding, and a unified instruction tuning mechanism, making it particularly effective for handling complex remote sensing data. In addition to these novel approaches, many remote sensing VLMs build upon the foundational architecture of LLaVA, adapting it to better address the specific challenges of remote sensing. RS-LLaVA enhances the standard LLaVA framework by incorporating a low-rank adaptation (LoRA) mechanism, optimizing the model’s ability to process the multi-scale and high-resolution nature of remote sensing images while maintaining computational efficiency. Similarly, H2RSVLM extends LLaVA by integrating a pretrained vision encoder and LLM, connected through a multilayer perceptron (MLP) projector. This model is particularly designed to improve its handling of complex queries and mitigate incorrect generations, especially in remote sensing contexts where ”hallucinations” are a common issue. GeoChat, also based on LLaVA, incorporates task-specific prompts and spatial capabilities, enabling it to handle multiple remote sensing tasks more effectively. These enhancements allow the model to better understand visual prompts and generate grounded object descriptions, providing more detailed and context-specific interactions in response to image and region-level queries.
@@ III-B Model Architecture
As shown in Figure 1, the proposed RS-MoE model consists of four modules, including an Image Encoder, a Vision Language Model (VLM) Encoder, a Large Language Model (LLM) Block, and a Mixture of Experts (MoE) Block.
@@ III-B1 Image Encoder
Given a remote sensing image $I$, we first employ an Image Encoder $f_{I}$ to extract the visual feature $F_{I}$. We adopt the ViT-G/14 [24], a large Vision Transformer that contains nearly two billion parameters, as our image encoder. We freeze this visual backbone during the entire model training process.
@@ III-B2 VLM Encoder
Inspired by the InstructBLIP [13], our VLM Encoder, denoted as $f_{\text{VLM}}$, is specifically designed to extract visual features that are attuned to the given instructions $T$. Firstly, a series of learnable query embeddings, along with corresponding instructions, are fed into the self-attention layers of the VLM Encoder, yielding an intermediate feature set $F_{\text{SA}}$. Subsequently, these features, together with the visual features $F_{I}$ derived from the Image Encoder, are processed through cross-attention layers. This step is pivotal for synthesizing the instruction-aware visual features $F_{\text{CA}}$, which are tailored to be both adaptable and insightful based on the input instructions. Finally, the feature set $F_{\text{CA}}$ progresses through a feed-forward network, culminating in a fully connected layer that outputs the refined feature representation $F_{\text{VLM}}$.
@@ III-B3 LLM Block
The learned instruction-aware features $F_{\text{VLM}}$ and the corresponding instructions $T$ serve as input prompts to the LLM Block $f_{l}$, facilitating the generation of detailed and coherent natural language captions $S$. This LLM Block is used during the first training stage of our model to set up and initialize the weights of the LLM models within the MoE Block for next training stage.
@@ III-B4 MoE Block
Unlike traditional MoE modules, the MoE Block comprises a novel instruction-router, the MoE Block comprises an Instruction-router $R$ and $N$ distinct Large Language Models (LLMs) , denoted as $\{f_{l1},f_{l2},\ldots,f_{lN}\}$. The instruction-router $R$ is designed to generate $N$ detailed prompts based on the input instruction $T$, represented as $\{P_{1},P_{2},\ldots,P_{N}\}$. Its design is grounded in the principles of Prompt Learning, adapting to various characteristics of the input image and dynamically generating prompts related to visual features to guide the expert models in producing targeted descriptions. During training, it learns how to dynamically adjust prompts based on visual features and task instructions, thereby more effectively guiding each expert model’s generation process. This enables the model to adapt to different image features and generate more precise and detailed descriptions.
The prompts generated by the instruction-router are then directed to their respective LLMs, each with the precise objective of focusing the LLM on a specialized sub-task. This targeted approach enables each LLM to concentrate on a specific aspect of the input data, allowing the model to develop a deep, focused expertise within a defined domain. Specifically, we break down the caption generation task into three specialized subtasks, assigning each LLM to one of the following areas: thematic interpretation, object identification, and the analysis of object relationships. By distributing these distinct sub-tasks to separate LLMs, the model effectively reduces the complexity of the captioning task into more manageable components, allowing each LLM to excel in solving its designated aspect of the task. This modular design enhances precision, as each LLM specializes in a narrower scope, reducing the risk of generalization errors. Additionally, this specialization allows for more nuanced control over the model’s outputs. The instruction-router dynamically adjusts prompts based on the context, ensuring that each LLM is aligned with the most relevant and accurate information for its sub-task. Ultimately, this framework enhances the model’s overall performance by leveraging the collective expertise of multiple LLMs, each optimized for a particular sub-task in the caption generation process.

### v2

# RS-MoE: A Vision-Language Model with Mixture of Experts for Remote Sensing Image Captioning and Visual Question Answering

@@ Abstract
Remote Sensing Image Captioning (RSIC) presents unique challenges and plays a critical role in applications such as environmental monitoring, urban planning, and disaster management. Traditional RSIC methods often struggle to produce rich and diverse descriptions. Recently, with significant advancements in Vision-Language Models (VLMs), efforts have emerged to integrate these models into the remote sensing domain and to introduce richly descriptive datasets specifically designed to enhance VLM training. However, most current RSIC models generally apply only fine-tuning to these datasets without developing models tailored to the unique characteristics of remote sensing imagery. This paper proposes RS-MoE, the first Mixture of Expert based VLM specifically customized for remote sensing domain. Unlike traditional MoE models, the core of RS-MoE is the MoE Block, which incorporates a novel Instruction Router and multiple lightweight Large Language Models (LLMs) as expert models. The Instruction Router is designed to generate specific prompts tailored for each corresponding LLM, guiding them to focus on distinct aspects of the RSIC task. This design not only allows each expert LLM to concentrate on a specific subset of the task, thereby enhancing the specificity and accuracy of the generated captions, but also improves the scalability of the model by facilitating parallel processing of sub-tasks. Additionally, we present a two-stage training strategy for tuning our RS-MoE model to prevent performance degradation due to sparsity. We fine-tuned our model on the RSICap dataset using our proposed training strategy. Experimental results on the RSICap dataset, along with evaluations on other traditional datasets where no additional fine-tuning was applied, demonstrate that our model achieves state-of-the-art performance in generating precise and contextually relevant captions. Notably, our RS-MoE-1B variant achieves performance comparable to 13B VLMs, demonstrating the efficiency of our model design. Moreover, our model demonstrates promising generalization capabilities by consistently achieving state-of-the-art performance on the Remote Sensing Visual Question Answering (RSVQA) task.
@@ I Introduction
Image captioning integrates computer vision and natural language processing to automatically generate descriptive text for images, effectively bridging the information gap between visual and linguistic domains. Unlike captioning of common natural images, remote sensing image captioning (RSIC) poses greater challenges due to the inherent characteristics of remote sensing images, which cover larger geographic areas and more diverse geographical objects [1, 2]. Consequently, RSIC tasks demand not only the description of geographic object information but also a comprehensive understanding of relationships and interactions among various geographic objects.
Early studies of RSIC focused on traditional methodologies, including template-based  [3] and retrieval-based methods [4, 5]. However, these methods cannot generate rich and varied descriptive sentences. Recent studies [6, 7, 8, 9, 10, 11] commonly adopt the encoder-decoder architecture, which divides the RSIC task into an image encoding phase that extracts semantic features from the input image, and a sequence modeling phase that uses the extracted features to generate text and sentences. Depending on the specific models employed, encoder-decoder approaches can be further categorized into CNN-based encoders with RNN/LSTM decoders [7], and CNN-based encoders with Transformer decoders [8]. Although these methods have achieved satisfying performance on the RSIC task, they typically generate only one or two simple sentences for captioning, limiting their practical application. This limitation typically arises due to two principal reasons: the simplicity and repetitiveness of the sentences in the datasets used for training models; and the relatively limited ability of these models to extract semantic features and generate complex descriptions.
The Large Language Models (LLMs) and Vision Language Models (VLMs) have recently achieved significant success across multiple fields, including computer vision [12], natural language processing [13], and robotics [14, 15]. Particularly, VLMs [16, 17, 18] have effectively narrowed the gap between visual images and textual natural language by advancing the understanding of intermodal relationships, reaching a level of visual comprehension comparable to human capabilities. Therefore, some researchers [19, 20, 21, 22, 23, 24, 25, 26] have recently shifted their focus to applying VLMs to remote sensing visual interpretation tasks. Specifically, for RSIC tasks, the RSGPT [20] introduced the human-annotated RSICap dataset, which provides high-quality and detailed descriptions for remote sensing images. The remote sensing captions in this dataset include summarized scene descriptions, object information (color, shape, and count), and object relationships (relative position). However, RSGPT merely fine-tunes an existing VLM model on this dataset without proposing a novel model to explore and utilize the provided detailed descriptions deeply.
@@
  • We are the first to introduce the MoE framework to multimodal remote sensing, leveraging its task decomposition concept with specialized expert models to effectively address the complexity and diversity of remote sensing data across both visual and textual modalities.
  • We propose RS-MoE, a novel Vision-Language Model specifically designed for remote sensing image captioning, with an Instruction Router that dynamically generates task-specific prompts and lightweight LLMs as expert models to enhance both effectiveness and efficiency.
  • We present a two-stage training strategy for RS-MoE, incorporating proper initialization to mitigate sparsity-induced degradation and employing LoRA to reduce trainable parameters, improving both efficiency and manageability of the model during the training process.
  • Extensive experiments demonstrate that, with fine-tuning on a single dataset $\sim 3,000$ images, our model achieves state-of-the-art performance on five RSIC datasets and strong generalization on two RSVQA datasets. Notably, the lightweight RS-MoE-1B matches the performance of larger 13B VLMs while being significantly more efficient.
Figure 1: Overview of the proposed RS-MoE model, which consists of four key components: the Image Encoder, the VLM Encoder, the LLM Block and the MoE Block. The MoE Block comprises an Instruction Router that dynamically generates task-specific prompts and three lightweight LLMs, which focus on different aspects of the captioning task. In the generated captions, shown in the top right corner of the figure, distinct colors represent each aspect: orange for the overall theme, purple for specific objects, and green for relationships between objects. RS-MoE is trained using a novel two-stage training strategy specifically designed for remote sensing image captioning. In Stage I (a), the VLM Encoder and the LLM Block are fine-tuned to initialize model weights specifically designed for the RSIC task. In Stage II (b), the MoE Block is fine-tuned to produce more detailed captions for RSIC tasks.
@@ II-B VLM in Remote Sensing
Vision-Language Models (VLMs) [16, 17, 18] are advanced frameworks that integrate visual and textual modalities, enabling the interpretation and generation of visual content with corresponding linguistic descriptions. These models have achieved remarkable progress in tasks involving natural images, particularly in generating dialogues and interpreting visual data. By combining visual features with contextual cues, VLMs offer an interactive and coherent understanding of visual inputs [27, 28].
@@
Inspired by the rapid advancements in computer vision, researchers have increasingly focused on integrating Vision-Language Models (VLMs) into remote sensing tasks. One of the early comprehensive reviews [29], highlighted key VLM applications within the remote sensing domain, identifying both the current challenges and future directions for research. Since then, several VLM-based models have emerged to address these challenges [19, 20, 21, 22, 23, 24, 25, 30]. For example,  [23] introduced EarthGPT, a model specifically designed for multi-sensor remote sensing image interpretation. This model integrates three core innovations: vision-enhanced perception, cross-modal understanding, and a unified instruction tuning mechanism, making it particularly effective for handling complex remote sensing data. In addition to these novel approaches, many remote sensing VLMs build upon the foundational architecture of LLaVA, adapting it to better address the specific challenges of remote sensing. RS-LLaVA enhances the standard LLaVA framework by incorporating a low-rank adaptation (LoRA) mechanism, optimizing the model’s ability to process the multi-scale and high-resolution nature of remote sensing images while maintaining computational efficiency. Similarly, H2RSVLM [22] extends LLaVA by integrating a pretrained vision encoder and LLM, connected through a multilayer perceptron (MLP) projector. This model is particularly designed to improve its handling of complex queries and mitigate incorrect generations, especially in remote sensing contexts where ”hallucinations” are a common issue. GeoChat [21], also based on LLaVA, incorporates task-specific prompts and spatial capabilities, enabling it to handle multiple remote sensing tasks more effectively. These enhancements allow the model to better understand visual prompts and generate grounded object descriptions, providing more detailed and context-specific interactions in response to image and region-level queries.
II-C Mixture of Experts
The Mixture of Experts (MoE) framework, first introduced by [31], is a foundational technique in machine learning designed to dynamically allocate computational resources by selectively activating specialized experts for specific tasks. This approach has been extensively studied across diverse domains, including recommender systems, natural language processing, computer vision, and multimodal learning. For example, in recommender systems, the Multi-gate Mixture of Experts (MMoE) [32] was developed as a multi-task learning model that explicitly models task relationships by sharing expert submodels across tasks, with separate gating networks optimizing individual tasks. Similarly, in natural language processing, the Stable Transferable Mixture of Experts (ST-MoE) [33] enhances the practicality and robustness of sparse models, achieving state-of-the-art performance on a wide range of natural language benchmarks. In computer vision, the Vision Mixture of Experts (V-MoE) [34] introduced a sparsely activated transformer model capable of scaling to 15 billion parameters, setting a new benchmark for efficiency and scalability in vision tasks. In the multimodal domain, the Language-Image Mixture of Experts (LIMoE) [35] integrates vision and language modalities through contrastive training, enabling expert layers to dynamically partition data based on modalities. Furthermore, the MoE-LLaVA [36] activates only the top-k experts during inference, optimizing computational efficiency for large vision-language models.
In contrast, the application of MoE in remote sensing remains largely underexplored. For instance, the MV-MoE [37] leverages a multi-task learning framework by incorporating attention mechanisms to represent multi-modal images across various scenes adaptively. Similarly, the SW-MoE [38] employs an MoE structure to activate experts based on different image scales and scenes, addressing challenges such as complex backgrounds and imbalanced datasets. Despite these advancements, existing studies are confined to single-modality applications. To the best of our knowledge, no prior work has explored the use of MoE in remote sensing for multimodal tasks. In this work, we address this gap by proposing a novel MoE-based VLM specifically designed to tackle the unique challenges of remote sensing image captioning and VQA.
@@ III-B Model Architecture
As shown in Fig. 1, the proposed RS-MoE model consists of four modules, including an Image Encoder, a Vision Language Model (VLM) Encoder, a Large Language Model (LLM) Block, and a Mixture of Experts (MoE) Block.
@@ III-B1 Image Encoder
Given the remote sensing image $I$ as defined earlier, we employ an Image Encoder $f_{I}$ to extract its visual features $F_{I}$. Specifically, $F_{I}=f_{I}(I)$, where $F_{I}\in\mathbb{R}^{C\times W^{\prime}\times H^{\prime}}$, $C$ is the number of feature channels, and $W^{\prime}$ and $H^{\prime}$ represent the downsampled spatial dimensions. For this purpose, we adopt ViT-G/14 [39], a Vision Transformer with nearly two billion parameters, as the backbone of the Image Encoder. To prevent overfitting and ensure stable feature extraction, the parameters of $f_{I}$ are frozen during the entire training process.
@@ III-B2 VLM Encoder
Inspired by InstructBLIP [18], our VLM Encoder, denoted as $f_{\text{VLM}}$, is specifically designed to generate visual features aligned with the given instructions $T$. The process begins with a set of learnable query embeddings $Q_{T}\in\mathbb{R}^{L\times C^{\prime}}$, where $L$ is the number of queries and $C^{\prime}$ is the feature dimension. These queries, combined with the instructions $T$, are processed through the self-attention layers of the encoder to produce an intermediate instruction-aware feature set $F_{\text{SA}}$, defined as:
$F_{\text{SA}}=\textit{SelfAttention}(Q_{T},T)$ (1)
Next, the features $F_{\text{SA}}$ are integrated with the visual features $F_{I}$ extracted by the Image Encoder using cross-attention layers, resulting in instruction-aware visual features $F_{\text{CA}}$:
$F_{\text{CA}}=\textit{CrossAttention}(F_{\text{SA}},F_{I})$ (2)
This step dynamically aligns the visual features with the task-specific instructions, tailoring them to the input context. Finally, the refined features $F_{\text{CA}}$ pass through a feed-forward network and a fully connected layer to yield the final feature representation $F_{\text{VLM}}$:
$F_{\text{VLM}}=\text{FFN}(F_{\text{CA}})$ (3)
By incorporating these steps, the VLM Encoder ensures that the extracted features are both adaptable and insightful, effectively aligning visual and instruction-specific information to support downstream tasks.
@@ III-B3 LLM Block
In the first training stage, The learned instruction-aware features $F_{\text{VLM}}$, along with the corresponding instructions $T$, serve as input prompts to the LLM Block $f_{l}$. The LLM Block generates a descriptive caption $S=\{s_{1},\cdots,s_{n}\}$ that effectively summarizes the content of the remote sensing image $I$ by identifying the objects in $\mathcal{O}$ and interpreting their relationships in $\mathcal{R}$. Formally, this process can be expressed as:
$S=f_{l}(F_{\text{VLM}},T)$ (4)
In the first training stage, the LLM Block is fine-tuned to initialize the weights of the LLM models used in the MoE Block. This preparatory step ensures that the models are capable of accurately capturing both object-level details and relationship dynamics, equipping the system to generate precise and coherent captions in subsequent stages.
@@ III-B4 MoE Block
Similarly, in the second training stage, the features generated by the VLM Encoder, $F_{\text{VLM}}$, are directly input into the MoE Block to facilitate caption generation. Unlike traditional MoE modules, the MoE Block comprises a novel instruction-router $R$ and $N$ distinct LLMs, denoted as $\{f_{l1},f_{l2},\ldots,f_{lN}\}$. The instruction-router $R$ dynamically generates $N$ detailed and adaptive prompts based on the input instruction $T$ and the instruction-aware VLM features $F_{\text{VLM}}$:
$\{P_{1},P_{2},\ldots,P_{N}\}=R(T,F_{\text{VLM}})$ (5)
Here, $R$ incorporates a feed-forward network (FFN) that projects the combined instruction $T$ and visual features $F_{\text{VLM}}$ into the embedding space for each sub-task. This dynamic prompt generation ensures that each $P_{i}$ is tailored to its respective sub-task. These tailored prompts are then directed by the instruction router to their corresponding LLMs, enabling each LLM to focus on producing precise and contextually relevant outputs for its assigned sub-task. In our experiment, the task is divided into three subtasks: thematic interpretation, object identification, and analysis of object relationships. Each LLM $f_{li}$ focuses on one sub-task, with its output expressed as:
$s_{i}=f_{li}(P_{i},F_{\text{VLM}}),\quad i=1,2,3$ (6)
The final caption $S$ is obtained by aggregating the outputs of all LLMs:
$S=\text{Aggregate}(\{s_{1},s_{2},s_{3}\})$ (7)
By minimizing the loss of each LLM output against its respective subtask-specific ground truth, each LLM is guided to specialize in a narrow aspect of the task, reducing complexity and enhancing precision. This modular design, combined with dynamically adjusted prompts during training, ensures that the instruction router provides task-relevant guidance. Generalization errors are computed based on the subtask-specific ground truth, ensuring that the evaluation aligns with each subtask’s intended learning objective. This approach prevents misleading assessments that could arise from evaluating outputs against an overly broad reference. Ultimately, this framework leverages the combined expertise of multiple LLMs to significantly improve caption generation performance.